# Association of the IceCube neutrinos with CAZ blazar light curves

### v1

@@ 4.6 Comparison to K24
The spatio-temporal results of this work are not directly comparable to those of K24 due to differences in how the temporal associations are defined. First, here we focus on rapid changes in the optical flux density, whereas in K24 we focused on the long-term, seasonal changes, allowing us to compare the optical behavior to the generally longer term radio behavior. Second, here the spatio-temporal tests are only applied to confidently variable sources, while in K24 we did not distinguish between variable or non-variable sources. Although the effect of variability in practice should be comparable in both studies, because non-variable sources had minimal effect on the tests of K24 due to their lack of major flares. Third, the more rapid temporal metrics of this study are more strongly affected by the presence of seasonal gaps (see Sect. 4.1) than the longer-term metrics of K24. As a result, there are several spatio-temporal associations in K24 which are identified only as spatial associations in this study.
@@
For example, the most notable K24 association missing here is of the LBL CAZJ0502+1338 (4FGL J0502.5+1340) with IC151114A (=). Notably, this blazar has a second neutrino association (IC190712A; =) and was associated with a neutrino hotspot (Buson et al. 2023). As seen in Fig. 7 of K24, IC151114A arrived at a historically high radio state and a relatively high optical state. However, around this time, the optical brightness is not high enough to be a BB95 as it only reaches the 93th percentile, and there is not enough cadence to identify a potential BBHOP flare. Likewise, the second neutrino (IC190712A) arrived close to an optical flare but in a seasonal gap. Therefore, even though CAZJ0502+1338 is not identified to be spatio-temporally neutrino-associated in this study, it is still a promising neutrino-emitting candidate.
@@
Despite such individual differences, one can still broadly compare the correlation strengths of this study to those of K24. In the most comparable scenario, here we have (for all RFC† via BB95) while we had in K24 (for optical-only, counted TS in the enlarged scenario). The decrease in correlation strength is interesting because the radio-selected blazar sample of K24 (i.e., CGRaBS) contained only 1157 sources while the current radio-selected sample (i.e., RFC†) has 3225 sources (nearly three times as many sources and nearly two times as many optically variable sources). If a spatio-temporal correlation exists between neutrinos and the whole blazar population, increasing the size of the blazar sample should strengthen the correlation (e.g., Liodakis et al. 2022a). Thus, this decrease in the correlation strength is evidence against a general spatio-temporal connection between blazars and neutrinos, at least in the optical band. Crucially, this does not exclude a potential spatio-temporal connection in other bands than optical. The challenge is that in other bands it is not possible to check the effect of increasing the number of blazar light curves as done for optical here. For example, in H21 and K24, we already tested the largest possible number of long-term light curves currently available in the radio band. Nonetheless, the lack of a population-based blazar-neutrino correlation suggested by our optical analyses is in agreement with recent population studies (e.g., Abbasi et al. 2025) and generally inline with the estimate that only a small fraction IceCube neutrinos originate from blazars (e.g., Oikonomou 2022; Yoshida et al. 2023; Robinson & Böttcher 2024; also see Sect. 4.4).
@@ 4.7 Physical properties of the neutrino-associated blazars
Here we investigate if the physical properties of the neutrino-associated blazars are different from the rest of the blazar population. As discussed in Sect. 4.1, we may be missing 22% of potential spatio-temporal associations due to seasonal gaps. Therefore, in this section we consider all spatially neutrino-associated blazars.
@@
We select neutrino-associated blazars as those which are spatially associated with high-weight neutrinos (; see Sect. 4.5 for justification) regardless of their variability. While many of these are expected to be false candidates, at least we ensure that any spatially neutrino-associated blazar that could also potentially be temporally associated is included in the comparison. This approach has the added benefit of allowing for some meaningful statistics due to a larger number of associations, which is not possible when only focusing on the few high-weight spatio-temporal associations. We find that a total of 61 blazars from either the RFC† or the 4LAC samples are spatially associated with these 79 high-weight events. In terms of classification, 4 (7% of 61) are AGN, 16 (26%) unclassified blazars, 16 (26%) BLLs, and 25 (41%) FSRQs.
@@
In Fig. 7 we plot the distribution of four physical properties (synchrotron peak frequency, redshift, radio variability Doppler factor, and median X-ray flux density) and two light curve properties (CRTS and ZTF ) of all 5880 blazars of the RFC† and 4LAC samples (using blue, solid lines; denoted as “a”) along with spatially neutrino-associated blazars which have (using red, dash-dotted lines; denoted as “b”). To check if these distributions are significantly different, we use the Kolmogorov-Smirnov (KS) and Anderson-Darling (AD) tests. Although, for brevity, we only report the results of the KS test, which leads to more conservative correlation strengths in all scenarios. Since the sample sizes between the (a) and (b) distributions are vastly different (two orders of magnitude), we also apply the KS and AD tests to a randomly selected subsample of the larger (a) which is equal in size to the smaller (b). The resampling process is repeated times. We then calculate the fraction of the resampled KS and AD -values that cross the 2 threshold for each of the six physical quantities. By chance, only 4.6% of the resampled -values should cross the 2 threshold: a larger fraction would suggest that (a) and (b) are likely different distributions. For comparison, we also plot the respective distributions of the spatio-temporally neutrino-associated blazars via BB95 with , which were discussed in detail in Sect. 4.5. These are labeled as “c” and plotted using black, dotted lines. Unfortunately, the sample sizes of (c) are too small to allow for reliable KS or AD testing (even via resampling).
@@
Figure 7 (i) shows that spatially neutrino-associated blazars have similar average synchrotron peak frequencies with respect to other blazars in the RFC† and 4LAC samples. Meanwhile, the four high-weight, spatio-temporally neutrino-associated blazars (including CAZJ0211+1051 and CAZJ0509+0541) clearly show an excess around the ISP regime . This suggests that, while a few IBLs could be neutrino emitters, the entire IBL subclass may not be more efficient at emitting neutrinos than other blazar subclasses.
@@
As evident from Fig. 7 (ii), the neutrino-associated blazars seem to have a similar redshift distribution as the RFC† and 4LAC blazars. This is generally inline with the results of Azzollini et al. (2025) who studied the physical properties of 52 blazars which were spatially correlated with IceCube neutrino hotspots. This suggests that blazars at any redshift epoch could be neutrino emitters.
@@
Figure 7 (iii) hints that spatially neutrino-associated blazars have a slightly higher than average radio variability Doppler factor than the rest of the RFC† and 4LAC blazars. Although the significance of the simple KS test is less than 2, that of the simple AD test is 2 and the resampled KS test supports this implication (i.e., ) as does the resampled AD test. This is in agreement with the results of Plavin et al. (2025) who found that spatially neutrino-associated blazars have a higher than average Doppler factor with a significance of 2.4. To compare the results more directly, we repeat this test by selecting only gold events with , exactly as done by Plavin et al. (2025). In this case, the KS test gives 2.2 and reaches 28%, reaffirming the agreement. Remarkably, this agreement persists even though we have a total of 858 Doppler factors while Plavin et al. (2025) only had 205. Moreover, we have 13 neutrino-associated blazars with a Doppler factor estimate while they had 8. This apparent positive correlation between the Doppler factor and neutrino emission supports the idea that neutrinos are produced in the jet, such that a more highly beamed jet results in a higher neutrino flux (e.g., Kovalev et al. 2025). Nevertheless, the two high-weight, spatio-temporally neutrino-associated blazars with Doppler factor estimates are not highly beamed ( of 8.4 and 14.7).
@@
Figure 7 (iv) suggests that the spatially neutrino-associated blazars are on average slightly brighter in X-rays ( keV) than the rest of the RFC† and 4LAC blazars. Although the KS significance is just below 2, that of the AD test is 2 and the resampled KS test supports this (i.e., ) as does the resampled AD test. Additionally, the four high-weight, spatio-temporally neutrino-associated blazars clearly have an above-average X-ray flux density. This generally agrees with the results of Plavin et al. (2024) who found a hint of a spatial correlation between hard-X-ray-detected (10 keV) blazars and IceCube neutrinos at least at the 2 level. These results would suggest that X-ray-bright blazars are more efficient at producing IceCube neutrinos, which is inline with the theory of neutrino production. This process is most efficient at producing PeV neutrinos when the seed photons have keV (X-ray) energies in the proton frame (e.g., Roulet & Vissani 2021) from a dense ambient medium (e.g., Reimer et al. 2019), the synchrotron emission of the protons themselves (e.g., Mastichiadis & Petropoulou 2021), or the synchrotron or inverse-Compton emission of cospatial electrons (e.g., Cerruti et al. 2019). Nevertheless, we caution that this result is not obtained when we run more rigorous correlation tests on the whole X-ray-detected and X-ray-complete samples of blazars in our companion paper Paggi et al. (2025).
@@
Figure 7 (v) and (vi) show that the optical variability of the neutrino-emitting blazar candidates is similar to the rest of the RFC† and 4LAC blazars, although the four high-weight spatio-temporally neutrino-associated blazars have either above average CRTS or ZTF . This is due to a selection bias, since we only identify blazar flares in confidently variable light curves, which have critically large CRTS or ZTF by construction (for more details, see K25-Cat).
@@ 5 Summary and conclusions
We test the spatio-temporal correlation of 356 high-energy neutrino events of the updated IceCat1+ sample with 3225 and 3814 blazars of the radio-selected RFC† and -ray-selected 4LAC samples, respectively, together constituting 5880 unique blazars. We use the methodology of our previous studies, H21 and K24, optimized in K25-Sim. We utilize optical all-sky survey light curves, collected in K25-Cat, for as many blazars as possible. In short, we calculate the weighted sum of the number of neutrinos whose arrival time is coincident with a flare (defined in K25-Cat) in the light curve of a spatially associated blazar. The correlation strength is then obtained by comparing this weighted sum to those arising from randomized simulations. Lastly, we also perform a complementary analysis using only spatially neutrino-associated blazars. We summarize our conclusions below:
Despite a substantial increase in data, the number of confident spatio-temporal associations still remains relatively low, resulting in generally weak correlation strengths. However, the number of spatio-temporal associations could be underestimated by 22% due to the prevalence of seasonal gaps in the CAZ light curves.
The spatio-temporal association of the IBLs CAZ J0211+1051 and CAZJ0509+0541 with the two reliably reconstructed neutrino events IC131014A and IC170922A, respectively, dominate our correlation strengths. As a result, in only one of the 18 tests, the post-trial spatio-temporal correlation strength exceeds 2 (i.e., 2.61 between neutrinos and BB95 at the peak of a BBHOP flare in RFC† BLLs).
The strongest spatio-temporal correlations are obtained at the time of the peak of the most prominent optical flares.
We estimate an upper limit of 8% for the fraction cosmic neutrinos emitted during major optical flares from blazars, which is in agreement with previous estimates.
@@
Spatially neutrino-associated blazars have a slightly higher than average X-ray brightness, inline with the need for X-ray seed photons to produce neutrinos in the detectable range of IceCube. However, more rigorous tests done on X-ray-selected blazars in our companion paper Paggi et al. (2025) do not support this.
Spatially neutrino-associated blazars seem to have similar synchrotron peak frequency, redshift, and optical variability distributions as the general blazar population.
@@
Our spatio-temporal methodology, in combination with multiwavelength spectro-polarimetric observations, can be applied to data from the next-generation neutrino and electromagnetic facilities to elucidate the blazar-neutrino mystery (e.g., by testing predictions made for time delays between neutrino emission and electromagnetic flares; Podlesnyi & Oikonomou 2025). This is especially true for the southern sky owing to such upcoming facilities as KM3NeT (e.g., Adrián-Martínez et al. 2016), Cherenkov Telescope Array Observatory (CTAO; Cherenkov Telescope Array Consortium et al. 2019), Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; e.g., Ivezić et al. 2019), Square Kilometre Array Observatory (SKAO; e.g., Agudo et al. 2015), and the African Milimetre Telescope (AMT; Backes et al. 2016).

### v2

@@ 4.6 Comparison to K24
The spatiotemporal results of this work are not directly comparable to those of K24 due to differences in how the temporal associations are defined. First, here we focus on rapid changes in the optical flux density, whereas in K24 we focused on the long-term, seasonal changes, allowing us to compare the optical behavior to the generally longer term radio behavior. Second, here the spatiotemporal tests are only applied to confidently variable sources, while in K24 we did not distinguish between variable or non-variable sources. Although the effect of variability in practice should be comparable in both studies, because non-variable sources had minimal effect on the tests of K24 due to their lack of major flares. Third, the more rapid temporal metrics of this study are more strongly affected by the presence of seasonal gaps (see Sect. 4.1) than the longer-term metrics of K24. As a result, there are several spatiotemporal associations in K24 which are identified only as spatial associations in this study.
@@
For example, the most notable K24 association missing here is of the LBL CAZJ0502+1338 (4FGL J0502.5+1340) with IC151114A (=). Notably, this blazar has a second neutrino association (IC190712A; =) and was associated with a neutrino hotspot (Buson et al. 2023). As seen in Fig. 7 of K24, IC151114A arrived at a historically high radio state and a relatively high optical state. However, around this time, the optical brightness is not high enough to be a BB95 as it only reaches the 93th percentile, and there is not enough cadence to identify a potential BBHOP flare. Likewise, the second neutrino (IC190712A) arrived close to an optical flare but in a seasonal gap. Therefore, even though CAZJ0502+1338 is not identified to be spatiotemporally neutrino-associated in this study, it is still a promising neutrino-emitting candidate.
@@
Despite such individual differences, one can still broadly compare the correlation strengths of this study to those of K24. In the most comparable scenario, here we have (for all RFC† via BB95) while we had in K24 (for optical-only, counted TS in the enlarged scenario). The decrease in correlation strength is interesting because the radio-selected blazar sample of K24 (i.e., CGRaBS) contained only 1157 sources while the current radio-selected sample (i.e., RFC†) has 3225 sources (nearly three times as many sources and nearly two times as many optically variable sources). If a spatiotemporal correlation exists between neutrinos and the whole blazar population, increasing the size of the blazar sample should strengthen the correlation (e.g., Liodakis et al. 2022a). Thus, this decrease in the correlation strength is evidence against a general spatiotemporal connection between blazars and neutrinos, at least in the optical band. Crucially, this does not exclude a potential spatiotemporal connection in other bands than optical. The challenge is that in other bands it is not possible to check the effect of increasing the number of blazar light curves as done for optical here. For example, in H21 and K24, we already tested the largest possible number of long-term light curves currently available in the radio band. Nonetheless, the lack of a population-based blazar-neutrino correlation suggested by our optical analyses is in agreement with recent population studies (e.g., Abbasi et al. 2025) and generally inline with the estimate that only a small fraction IceCube neutrinos originate from blazars (e.g., Oikonomou 2022; Yoshida et al. 2023; Robinson & Böttcher 2024; also see Sect. 4.4).
@@ 4.7 Physical properties of the neutrino-associated blazars
Here, we investigate if the physical properties of the neutrino-associated blazars are different from the rest of the blazar population. As discussed in Sect. 4.1, we may be missing 22% of potential spatiotemporal associations due to seasonal gaps. Therefore, in this section we consider all spatially neutrino-associated blazars.
@@
We selected neutrino-associated blazars on the basis of an evident spatial association with high-weight neutrinos (; see Sect. 4.5 for justification) regardless of their variability. While many of these are expected to be false candidates, at least we ensured that any spatially neutrino-associated blazar that could also potentially be temporally associated is included in the comparison. This approach has the added benefit of allowing for some meaningful statistics due to a larger number of associations, which is not possible when only focusing on the few high-weight spatiotemporal associations. We find that a total of 61 blazars from either the RFC† or the 4LAC samples are spatially associated with these 79 high-weight events. In terms of classification, 4 (7% of 61) are AGN, 16 (26%) unclassified blazars, 16 (26%) BLLs, and 25 (41%) FSRQs.
@@
In Fig. 7, we plot the distribution of four physical properties (synchrotron peak frequency, redshift, radio variability Doppler factor, and median X-ray flux density) and two light curve properties (CRTS and ZTF ) of all 5880 blazars of the RFC† and 4LAC samples (using blue, solid lines; denoted as “a”) along with spatially neutrino-associated blazars which have (using red, dash-dotted lines; denoted as “b”). To check if these distributions are significantly different, we use the Kolmogorov-Smirnov (KS) and Anderson-Darling (AD) tests. Although, for brevity, we only report the results of the KS test, which leads to more conservative correlation strengths in all scenarios. Since the sample sizes between the (a) and (b) distributions are vastly different (two orders of magnitude), we also apply the KS and AD tests to a randomly selected subsample of the larger (a) which is equal in size to the smaller (b). The resampling process is repeated times. We then calculate the fraction of the resampled KS and AD -values that cross the 2 threshold for each of the six physical quantities. By chance, only 4.6% of the resampled -values should cross the 2 threshold: a larger fraction would suggest that (a) and (b) are likely different distributions. For comparison, we also plot the respective distributions of the spatiotemporally neutrino-associated blazars via BB95 with , which were discussed in detail in Sect. 4.5. These are labeled as “c” and plotted using black, dotted lines. Unfortunately, the sample sizes of (c) are too small to allow for reliable KS or AD testing (even via resampling).
@@
Figure 7 (i) shows that spatially neutrino-associated blazars have similar average synchrotron peak frequencies with respect to other blazars in the RFC† and 4LAC samples. Meanwhile, the four high-weight, spatiotemporally neutrino-associated blazars (including CAZJ0211+1051 and CAZJ0509+0541) clearly show an excess around the ISP regime . Although the most likely neutrino-emitting blazars are IBLs, we do not find any evidence from the distribution of the spatially neutrino-associated blazars that the IBL subclass is more efficient at emitting neutrinos than other blazar subclasses.
@@
As made evident from Fig. 7 (ii), the neutrino-associated blazars seem to have a similar redshift distribution as the RFC† and 4LAC blazars. This is generally in line with the results of Azzollini et al. (2025) who studied the physical properties of 52 blazars which were spatially correlated with IceCube neutrino hotspots. This suggests that blazars at any redshift epoch could be neutrino emitters.
@@
Figure 7 (iii) hints that spatially neutrino-associated blazars have a slightly higher than average radio variability Doppler factor, , than the rest of the RFC† and 4LAC blazars. Although the significance of the simple KS test is less than 2, that of the simple AD test is 2 and the resampled KS test supports this implication (i.e., ) as does the resampled AD test. This is in agreement with the results of Plavin et al. (2025) who found that spatially neutrino-associated blazars have a higher than average Doppler factor with a significance of 2.4. To compare the results more directly, we repeat this test by selecting only gold events with , exactly as done by Plavin et al. (2025). In this case, the KS test gives 2.2 and reaches 28%, reaffirming the agreement. Remarkably, this agreement persists even though we have a total of 858 Doppler factors while Plavin et al. (2025) only had 205. Moreover, we have 13 neutrino-associated blazars with a Doppler factor estimate while they had 8. This apparent positive correlation between the Doppler factor and neutrino emission supports the idea that neutrinos are produced in the jet, such that a more highly beamed jet results in a higher neutrino flux (e.g., Kovalev et al. 2025). Nevertheless, the two high-weight, spatiotemporally neutrino-associated blazars with Doppler factor estimates are not highly beamed ( of 8.4 and 14.7).
@@
Figure 7 (iv) suggests that the spatially neutrino-associated blazars are on average slightly brighter in X-rays ( keV) than the rest of the RFC† and 4LAC blazars. Although the KS significance is just below 2, that of the AD test is 2 and the resampled KS test supports this (i.e., ) as does the resampled AD test. Additionally, the four high-weight, spatiotemporally neutrino-associated blazars clearly have an above-average X-ray flux density. This generally agrees with the results of Plavin et al. (2024) who found a hint of a spatial correlation between hard-X-ray-detected (10 keV) blazars and IceCube neutrinos at least at the 2 level. These results would suggest that X-ray-bright blazars are more efficient at producing IceCube neutrinos, which is inline with the theory of neutrino production. This process is most efficient at producing PeV neutrinos when the seed photons have keV (X-ray) energies in the proton frame (e.g., Roulet & Vissani 2021) from a dense ambient medium (e.g., Reimer et al. 2019), the synchrotron emission of the protons themselves (e.g., Mastichiadis & Petropoulou 2021), or the synchrotron or inverse-Compton emission of cospatial electrons (e.g., Cerruti et al. 2019). Nevertheless, we caution that this result is not obtained when we run more rigorous correlation tests on the whole X-ray-detected and X-ray-complete samples of blazars in our companion paper Paggi et al. (in prep.).
@@
Figure 7 (v) and (vi) show that the optical variability of the neutrino-emitting blazar candidates is similar to the rest of the RFC† and 4LAC blazars, although the four high-weight spatiotemporally neutrino-associated blazars have either above average CRTS or ZTF . This is due to a selection bias, since we only identify blazar flares in confidently variable light curves, which have critically large CRTS or ZTF by construction (for more details, see K26).
@@ 5 Summary and conclusions
We tested the spatiotemporal correlation of 356 high-energy neutrino events of the updated IceCat1+ sample with 3225 and 3814 blazars of the radio-selected RFC† and -ray-selected 4LAC samples, respectively, altogether constituting 5880 unique blazars. We used the methodology applied in our previous studies, H21 and K24, and optimized in K25. We utilized optical all-sky survey light curves, collected in K26, for as many blazars as possible. In short, we calculated the weighted sum of the number of neutrinos whose arrival time is coincident with a flare (defined in K26) in the light curve of a spatially associated blazar. The correlation strength was then obtained by comparing this weighted sum to those arising from randomized simulations. Lastly, we also performed a complementary analysis using only spatially neutrino-associated blazars. We summarize our conclusions below:
Despite a substantial increase in data, the number of confident spatiotemporal associations still remains relatively low, resulting in generally weak correlation strengths. However, the number of spatiotemporal associations could be underestimated by 22% due to the prevalence of seasonal gaps in the CAZ light curves.
The spatiotemporal association of the IBLs CAZ J0211+1051 and CAZJ0509+0541 with the two reliably reconstructed neutrino events IC131014A and IC170922A, respectively, dominate our correlation strengths. As a result, in only one of the 18 tests, the post-trial spatiotemporal correlation strength exceeds 2 (i.e., 2.61 between neutrinos and BB95 at the peak of a BBHOP flare in RFC† BLLs).
The strongest spatiotemporal correlations are obtained at the time of the peak of the most prominent optical flares.
We estimated an upper limit of 8% for the fraction of the updated IceCat1+ cosmic neutrinos emitted during major optical flares from blazars, which is in agreement with previous estimates.
@@
Spatially neutrino-associated blazars have a slightly higher than average X-ray brightness, in line with the need for X-ray seed photons to produce neutrinos in the detectable range of IceCube. However, more rigorous tests done on X-ray-selected blazars in our companion paper (Paggi et al., in prep.) do not support this.
Spatially neutrino-associated blazars seem to have similar synchrotron peak frequency, redshift, and optical variability distributions to those seen for the general blazar population.
@@
In the future, we aim to test the spatiotemporal correlation of the CAZ blazars against the upcoming IceCat-2 neutrinos (e.g., Zegarelli et al. 2025). IceCat-2 is expected to contain more events and, crucially, provide substantially improved spatial error regions as compared to IceCat-1. Notably, our spatiotemporal methodology, in combination with multiwavelength spectro-polarimetric observations, can be applied to data from the next-generation neutrino and electromagnetic facilities to elucidate the blazar-neutrino mystery (e.g., by testing predictions made for time delays between neutrino emission and electromagnetic flares; Podlesnyi & Oikonomou 2025). This is especially true for the southern sky owing to such upcoming facilities as KM3NeT (e.g., Adrián-Martínez et al. 2016), Cherenkov Telescope Array Observatory (CTAO; Cherenkov Telescope Array Consortium et al. 2019), Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; e.g., Ivezić et al. 2019), Square Kilometre Array Observatory (SKAO; e.g., Agudo et al. 2015), and the African Milimetre Telescope (AMT; Backes et al. 2016).
Data availability
Two tables, “neutrino.dat” and “assoc.dat”, are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/. The table “neutrino.dat” contains the full list of the updated IceCat1+ neutrinos. The table “assoc.dat” contains the full list of spatial and spatiotemporal blazar-neutrino associations. We note that Table 3 is a subset of “assoc.dat” in its entirety.